# Dynamics of Dehn Twists in the Outer Automorphism Group of a Free Group

### v1

@@ 3.2. Intersection number and Dehn twist
Conversely, assume that $\delta_{\vec{T}}$ and $\delta_{\vec{T}^{\prime}}$ commute. Then
$\delta_{\delta_{\vec{T}^{\prime}}(\vec{T})}=\delta_{\vec{T}^{\prime}}\delta_{\vec{T}}\delta_{\vec{T}^{\prime}}^{-1}=\delta_{\vec{T}}.$
So $\delta_{\vec{T}^{\prime}}(T)=T$. By Lemma 3.6, this implies that $i(T,T^{\prime})=0$. Therefore, $\delta_{\vec{T}}$ and $\delta_{\vec{T}^{\prime}}$ commute if and only if $i(T,T^{\prime})=0$. ∎
@@ 4. On the universal cover
Theorem 4.3 (Lyndon and Schützenberger [MR162838]).
@@ 5.4. Right-angled Artin group generated by powers of Dehn twists
In this section, we look into a specific collection of essential tori whose sufficiently large powers of Dehn twists generate a right-angled Artin group. Note that there exist finitely many Dehn twists that generate $\operatorname{Out}(\mathbb{F})$, which is more complicate than right-angled Artin groups. This section is inspired by Koberda [MR3000498] and Seo [MR4299673].
@@ 6. The space of factorial curves
Before proving Proposition 6.2, we require the following technical lemma. This result is analogous to [MR3053012, Lemma 4.2].
@@
The following proof is analogous to [MR3053012, Proposition 4.1].
@@
The following is motivated by [MR3053012, Proposition A.1].

### v2

@@ 3.2. Intersection number and Dehn twist
Conversely, assume that $\delta_{\vec{T}}$ and $\delta_{\vec{T}^{\prime}}$ commute. Then for each $k>0$, we have
$\delta_{\delta_{\vec{T}^{\prime}}^{k}(\vec{T})}=\delta_{\vec{T}^{\prime}}^{k}\delta_{\vec{T}}\delta_{\vec{T}^{\prime}}^{-k}=\delta_{\vec{T}}.$
So one gets $\delta_{\vec{T}^{\prime}}^{k}(T)=T$ for every $k>0$. By Lemma 3.6, this implies that $i(T,T^{\prime})=0$. Therefore, $\delta_{\vec{T}}$ and $\delta_{\vec{T}^{\prime}}$ commute if and only if $i(T,T^{\prime})=0$. ∎
@@ 4. On the universal cover
Theorem 4.3 (Lyndon and Schützenberger [16]).
@@ 5.4. Right-angled Artin group generated by powers of Dehn twists
In this section, we look into a specific collection of essential tori whose sufficiently large powers of Dehn twists generate a right-angled Artin group. Note that there exist finitely many Dehn twists that generate $\operatorname{Out}(\mathbb{F})$, which is more complicate than right-angled Artin groups. This section is inspired by Koberda [14] and Seo [17].
@@ 6. The space of factorial curves
Before proving Proposition 6.2, we require the following technical lemma. This result is analogous to [6, Lemma 4.2].
@@
The following proof is analogous to [6, Proposition 4.1].
@@
The following is motivated by [6, Proposition A.1].